VFTOVP(1) VFTOVP(1) NAME vftovp-convertvirtualfont(vf)filestovirtualpropertylists SYNOPSIS vftovp[-verbose][-charcode-format=format]vffilenametfmfilename[vplfilename] DESCRIPTION Thevftovpprogramtranslatesa(program-oriented)virtualfontfileanditsaccompanyingfont metricfiletoa(human-oriented)propertylistfile.Thus,thevirtualfontcanbeedited,andits exactcontentscanbedisplayedmnemonically.ItalsoservesasaVF-filevalidatingprogram;ifno errormessagesaregiven,theinputfilesarecorrect. Thevplfilename,mustbecomplete.Noaddingofdefaultextensionsisdone. ENVIRONMENT vftovplooksforvffilenamefirstinthecurrentdirectory.Ifitisnotpresentthere,itusesthe environmentvariableVFFONTS.Ifthatisnotset,itusesthevariableTEXFONTS.Ifthisis notset,itusesthesystemdefaults: VFFONTS:TEXFONTS, TEXFONTS:/usr/TeX/lib/tex/fonts//:., Seetex(1)forthedetailsofthesearching. OPTIONS Withoutthe-verboseoption,vftovpoperatessilently.Withit,abannerandprogressreportare printedonstdout. Theargumentformatto-charcode-formatspecifieshowcharactercodesareoutputinthePLfile. Bydefault,onlylettersanddigitsareoutputusingtheCintegercode(i.e.,inASCII);theothers areoutputinoctal.(Unlessthefont’scodingschemestartswithTeXmathsyorTeXmathex, thatis,whenallareoutputinoctal.)Ifformatisascii,allcharactercodesthatcorrespondto graphiccharacters,besidestheleftandrightparentheses,areoutputinASCII.Ontheother hand,ifformatisoctal,allcharactercodesareoutputinoctal. Seetex(1)forthedetailsofspecifyingoptions. SEEALSO vptovf(1) DonaldKnuth,VirtualFonts:MoreFunforGrandWizards,TUGboat11(1),April1990. AUTHORS DonaldKnuthwrotetheprogram,basedinpartonanideaofDavidFuchs,startingwiththecode fortftopl(1).KarlBerryadapteditforcompilationwithweb2c. 7Jan92 1